If you deposit $326 per quarter into an account which pays interest at a rate of 11% per year compounded quarterly, the amount of money you would have at the end of 23 years is equal to
Answer #1

Answer:

Number of Periods/years = 23

Nominal Rate of Interest = 11% = 0.11

Regular Withdrawal = 1,304

Benefit paid Annually

= 326 * 4 = 1304

T= (1 +

r = [1 + 0.11/4]^4 – 1 = 0.11462

(1 + r) -1 FV = A- r(P)

FV = 1304 * [(1 + 0.11462)^23 – 1]/ 0.11

= 131,959.7408

Thus the amount of money you would have at the end of 23 years is equal to $131,959.7408